1- Concord Corporation will receive $407000 in 7 years. If the appropriate interest rate is 8%, the present value of the $407000 receipt is?

2-Bonita Industries will receive $1400000 in 6 years. If the appropriate interest rate is 11%, the present value of the $1400000 receipt is?
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

1.Present value=407000/1.08^7

=407000*0.583490395

=$237480.59(Approx).

2.Present value=1400000/1.11^6

=1400000*0.534640836

=$748497.17(Approx).